# Tractable and Provably Efficient Distributional Reinforcement Learning with General Value Function Approximation

### v1

@@ 5 Algorithms
$\psi_{n}\Big{(}(\mathcal{B}_{r_{h^{\prime}}^{\tau}})_{\#}\bar{\eta}_{h}(s_{h^{\prime}+1}^{\tau})\Big{)}$ $=\mathbb{E}[(\bar{Z}_{h+1}^{k}(s_{h^{\prime}+1}^{\tau})+r_{h^{\prime}}^{\tau})^{n}]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}/H^{n-1}}$
$=\sum_{n^{\prime}=0}^{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H^{n^{\prime}}}\psi_{n^{\prime}}\Big{(}\bar{\eta}_{h}(s_{h^{\prime}+1}^{\tau})\Big{)}(r_{h^{\prime}}^{\tau})^{n-n^{\prime}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}/H^{n-1}}$
@@
$\tilde{f}^{k}_{h,\bar{\eta}}\leftarrow\arg\min_{f\in\mathcal{F}}\sum_{\tau=1}^{k-1}\sum_{h^{\prime}=1}^{H}\Big{(}\sum_{n=1}^{N}f^{(n)}(s_{h^{\prime}}^{\tau},a_{h^{\prime}}^{\tau})-\psi_{n}\Big{(}(\mathcal{B}_{r_{h^{\prime}}^{\tau}})_{\#}\bar{\eta}_{h+1}^{k}(s_{h^{\prime}+1}^{\tau})\Big{)}\Big{)}^{2}$
based on the dataset $\mathcal{D}^{k}_{h}$ which contains the sketch of temporal target $\psi_{1:N}\Big{(}(\mathcal{B}_{r_{h^{\prime}}^{\tau}})_{\#}\bar{\eta}_{h+1}^{k}(s_{h^{\prime}+1}^{\tau})\Big{)}$. Then we define $Q^{k}_{h}(\cdot,\cdot)=\min\{(\tilde{f}^{k}_{h,\bar{\eta}})^{(1)}(\cdot,\cdot)+b^{k}_{h}(\cdot,\cdot),H\}$ and choose the greedy policy $\pi^{k}_{h}(\cdot)$ with respect to $Q^{k}_{h}$. We repeat the procedure until all the $K$ episodes are completed.
Remark 5.1.
For an optimistic planning, we define the bonus function as the width function $b^{k}_{h}(s,a)\coloneqq w^{k}_{h}((\mathcal{F}^{N})^{k}_{h},s,a)$ where $(\mathcal{F}^{N})^{k}_{h}$ denotes a confidence region at step $h$, episode $k$. When $\mathcal{F}$ is a linear function class, the width function can be evaluated by simply computing the maximal distance of weight vector. For a general function class $\mathcal{F}$, computing the width function requires to solve a set-constrained optimization problem, which is known as NP-hard [17]. However, a width function is computed simply for optimistic exploration, and approximation errors are known to have a small effect on regret [1].
@@ 6 Theoretical Analysis
In this section, we provide the theoretical guarantees for SF-LSVI under assumption 4.8. We apply proof techniques from Wang et al. [50] and extend the result to a statistical functional lens. First, we generalize the concept of eluder dimension [43] to the vector-valued function, which has been widely used in RL literatures [4, 50, 27] to measure the complexity of learning with the function approximators.
@@
  • A state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ is $\bm{\epsilon}$-dependent on $\mathcal{Z}$ with respect to $\mathcal{F}^{N}$ if $\|f-g\|_{\mathcal{Z}}\leq\epsilon$ for any vector-valued function $f,g\in\mathcal{F}^{N}$, then $|f^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1)}}(s,a)-g^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1)}}(s,a)|\leq\epsilon$.
@@
We remark that using moment functional as a sketch is a unique structure that leads to a concentration results for the difference between the target and sampled sketches $\psi\Big{(}(\mathcal{B}_{r})_{\#}[\mathbb{P}\bar{\eta}]\Big{)}-\psi\Big{(}(\mathcal{B}_{r})_{\#}\bar{\eta}\Big{)}$, as the tranformation $\phi_{\psi}$ of moment functional is an identity. Furthermore, using normalized moments minimizes the size of the confidence region from $O(H^{N})$ to $O(\sqrt{N})$. This adjustment is akin to scaling the optimization function in multi-objective optimization to treat each objective equally, which effectively prevents the model from favoring objectives with larger scales.
@@
Under weaker structural assumptions, we show that SF-LSVI enjoys sublinear regret of order $\tilde{O}(d_{E}H^{\frac{3}{2}}\sqrt{K})$, which is $\sqrt{H}$ better than the state-of-the-art distributional framework algorithm V-EST-LSR [11]. For the linear MDP setting, we have $d_{E}=\tilde{O}(d)$ and thus the regret bound of SF-LSVI is $\tilde{O}(\sqrt{d^{2}H^{3}K})$ which matches a lower bound $\Omega(\sqrt{d^{2}H^{3}K})$ [53]. In our analysis, we highlight two main technical novelties which significantly reduces the complexity of regret;
@@
    We refine previous lemma of Osband et al. [39] and Wang et al. [50] to remove the dependency of $\beta(\mathcal{F}^{N},1/\delta)$ (See Appendix C.4), ensuring that regret bound depends only on the defined function class, not on the number of moment extracted.
@@ 7 Conclusions
We describe the sources of approximation error inherent in distribution-based updates and introduce a pivotal concept of Bellman unbiasedness, which enables to exactly learn the information of distribution. We also present a provably efficient distRL algorithm, SF-LSVI, with general value function approximation. Notably, our algorithm achieves a near-optimal regret bound of $\tilde{O}(d_{E}H^{\frac{3}{2}}\sqrt{K})$, matching the tightest upper bound achieved by non-distributional framework [53, 24]. One interesting future direction would be to reformulate the definition of regret as discrepencies in moments rather than the expected return, and to show the sample-efficiency of distRL. We hope that our work sheds some light on future research in analyzing the provable efficiency of distRL.
@@ Appendix A Notation
Appendix B Related Work and Discussion
B.1 Technical Clarifications on Linearity Assumption in Existing Results
@@
Rowland et al. [41] proved that quantile functional is not Bellman closed by providing a specific counterexample. However, their discussion based on counterexamples can be generalized as it assumes that the sketch Bellman operator for the quantile functional needs to be linear.
@@
Marthe et al. [35] proposed the notion of Bellman optimizable statistical functional which redefine the Bellman update by planning with respect to statistical functionals rather than expected returns. They proved that $W_{1}$-continuous Bellman Optimizable statistical functionals are characterized by exponential utilities $\frac{1}{\lambda}\log\mathbb{E}_{Z\sim\eta}[\exp(\lambda Z)]$. However, their proof requires some technical clarification regarding the assumption that such statistical functionals are linear.
@@
B.2 Existence of Nonlinear Bellman Closed Sketch.
@@
  • $\mathcal{T}_{\psi_{\text{max}}}\Big{(}\psi_{\text{max}}(\bar{\eta}(s))\Big{)}=\max_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\Big{(}\psi_{\text{max}}\big{(}(\mathcal{B}_{r})_{\#}\bar{\eta}(s^{\prime})\big{)}\Big{)}$
  • $\mathcal{T}_{\psi_{\text{min}}}\Big{(}\psi_{\text{min}}(\bar{\eta}(s))\Big{)}=\min_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\Big{(}\psi_{\text{min}}\big{(}(\mathcal{B}_{r})_{\#}\bar{\eta}(s^{\prime})\big{)}\Big{)}$.
B.3 Non-existence of sketch Bellman operator for quantile functional
@@
Definition B.1 (mixture-consistent).
A sketch $\psi$ is mixture-consistent if for any $\lambda\in[0,1]$ and any distributions $\eta_{1},\eta_{2}\in\mathscr{P}_{\psi}(\mathbb{R})$, there exists a corresponding function $h_{\psi}$ such that
$\psi(\lambda\eta_{1}+(1-\lambda)\eta_{2})=h_{\psi}\Big{(}\psi(\eta_{1}),\psi(\eta_{2}),\lambda\Big{)}.$
@@
$\mathbb{E}_{Z\sim\lambda\eta_{1}+(1-\lambda)\eta_{2}}[Z^{n}\exp(\lambda Z)]=\lambda\mathbb{E}_{Z\sim\eta_{1}}[Z^{n}\exp(\lambda Z)]+(1-\lambda)\mathbb{E}_{Z\sim\eta_{2}}[Z^{n}\exp(\lambda Z)].$
@@
Let $\lambda=\frac{1}{2}$ and $Z,Y$ be the random variables where $Z\sim\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{2}$ and $Y\sim\frac{1}{2}\delta_{k}+\frac{1}{2}\delta_{k+2}$. Then, $\text{Var}(Z)=\text{Var}(Y)=1$. While RHS is constant for any $k$, LHS is not a constant for any $k$, i.e.,
@@
$\text{Var}_{Z\sim\lambda\eta_{1}+(1-\lambda)\eta_{2}}[Z]$
$=\mathbb{E}_{Z\sim\lambda\eta_{1}+(1-\lambda)\eta_{2}}[Z^{2}]-(\mathbb{E}_{Z\sim\lambda\eta_{1}+(1-\lambda)\eta_{2}}[Z])^{2}$
$=\lambda\mathbb{E}_{Z\sim\eta_{1}}[Z^{2}]+(1-\lambda)\mathbb{E}_{Z\sim\eta_{2}}[Z^{2}]-(\lambda\mathbb{E}_{Z\sim\eta_{1}}[Z]+(1-\lambda)\mathbb{E}_{Z\sim\eta_{2}}[Z])^{2}$
$=\lambda(\text{Var}_{Z\sim\eta_{1}}[Z]+(\mathbb{E}_{Z\sim\eta_{1}}[Z])^{2})+(1-\lambda)(\text{Var}_{Z\sim\eta_{2}}[Z]+(\mathbb{E}_{Z\sim\eta_{2}}[Z])^{2})$
$\qquad-(\lambda\mathbb{E}_{Z\sim\eta_{1}}[Z]+(1-\lambda)\mathbb{E}_{Z\sim\eta_{2}}[Z])^{2}.$
@@
$\min_{Z\sim\lambda\eta_{1}+(1-\lambda)\eta_{2}}[Z]=\min(\min_{Z\sim\eta_{1}}[Z],\min_{Z\sim\eta_{2}}[Z])$
Since maximum and minimum functionals are mixture consistent, we can construct a nonlinear sketch bellman operator like the one in section B.2. This is possible because there is a nonlinear function $h_{\psi}$ that ensures the sketch is closed under mixture.
@@
Let $\lambda=\frac{1}{2}$ and $Z,Y$ be the random variables where $Z\sim 0.2\delta_{0}+0.8\delta_{1}$ and $Y\sim 0.6\delta_{0}+0.4\delta_{k}$ for some $10<k<$ 1. Then $\psi_{\text{med}}(Z)=1$ and $\psi_{\text{med}}(Y)=0$. However,
@@
For a given integer $N>0$ and a quantile level $\alpha\in[0,1]$, let $\lambda=\frac{1}{2}$ and a random variable $Y\sim p_{y_{0}}\delta_{0}+p_{y_{1}}\delta_{y_{1}}+\cdots+p_{y_{N}}\delta_{y_{N}}(0<y_{1}<\cdots<y_{N}<1)$ where $p_{y_{0}}>\alpha$ so that $\psi_{q_{\alpha}}(Y)=0$. Consider another random variable $Z\sim p_{z_{0}}\delta_{0}+p_{z_{1}}\delta_{1}$ where $p_{z_{0}}<\alpha$ so that $\psi_{q_{\alpha}}(Z)=1$. Then the $\alpha-$quantile of the mixture $X=\frac{Y+Z}{2}$ is
$\psi_{q_{\alpha}}[X]=y_{n}\text{ where }n=\min\left\{n\leq N\Big{|}\ \frac{1}{2}\sum_{n^{\prime}=0}^{n}p_{y_{n^{\prime}}}+\frac{1}{2}p_{z_{0}}>\alpha\right\}.$
Letting $p_{z_{0}}=2\alpha-\sum_{n^{\prime}=0}^{n}p_{y_{n^{\prime}}}$, we can manipulate $\psi_{q_{\alpha}}[X]$ to be any value of $y_{n}$. Hence, $\psi(q_{\alpha})[X]$ is a function of all possible outcomes of $Y$.
@@
Consider an MDP where initial state $s_{0}$ has no reward and transits to two state $s_{1},s_{2}$ with probability $\lambda,1-\lambda$ and reward distribution $\bar{\eta}_{1},\bar{\eta}_{2}$. Since $\psi$ is Bellman closed, $\psi(\bar{\eta}(s_{0}))$ is a function of $\psi(\bar{\eta}(s_{1}))$ and $\psi(\bar{\eta}(s_{2}))$, (i.e., $\psi(\bar{\eta}(s_{0}))$ = $g_{\psi}(\psi(\bar{\eta}(s_{1}))$, $\psi(\bar{\eta}(s_{2})))$ for some $g_{\psi}$). Since $\psi(\bar{\eta}(s_{0}))=\psi(\lambda\bar{\eta}(s_{1})+(1-\lambda)\bar{\eta}(s_{2}))$, it implies that $\psi$ is mixture-consistent. ∎
Combining the results of Lemma B.2 and Lemma B.3, we prove that a quantile sketch cannot be Bellman closed, no matter what additional sketches are provided.
Appendix C Proof

### v2

@@ 4 SF-LSVI: Statistical Functional Least Square Value Iteration
$\psi_{n}\Big{(}(\mathcal{B}_{r_{h^{\prime}}^{\tau}})_{\#}\bar{\eta}_{h}(s_{h^{\prime}+1}^{\tau})\Big{)}\coloneqq\frac{\mathbb{E}[(\bar{Z}_{h+1}^{k}(s_{h^{\prime}+1}^{\tau})+r_{h^{\prime}}^{\tau})^{n}]}{H^{n-1}}$
$\qquad\qquad\quad=\frac{\sum_{n^{\prime}=0}^{n}H^{n^{\prime}}\psi_{n^{\prime}}\Big{(}\bar{\eta}_{h}(s_{h^{\prime}+1}^{\tau})\Big{)}(r_{h^{\prime}}^{\tau})^{n-n^{\prime}}}{H^{n-1}}$
@@
$\tilde{f}^{k}_{h,\bar{\eta}}$ $\leftarrow\arg\min_{f\in\mathcal{F}}\sum_{\tau=1}^{k-1}\sum_{h^{\prime}=1}^{H}\Big{(}\sum_{n=1}^{N}f^{(n)}(s_{h^{\prime}}^{\tau},a_{h^{\prime}}^{\tau})$
$\qquad\qquad\qquad\qquad\quad-\psi_{n}\Big{(}(\mathcal{B}_{r_{h^{\prime}}^{\tau}})_{\#}\bar{\eta}_{h+1}^{k}(s_{h^{\prime}+1}^{\tau})\Big{)}\Big{)}^{2}$
based on the dataset $\mathcal{D}^{k}_{h}$ which contains the sketch of temporal target $\psi_{1:N}\Big{(}(\mathcal{B}_{r_{h^{\prime}}^{\tau}})_{\#}\bar{\eta}_{h+1}^{k}(s_{h^{\prime}+1}^{\tau})\Big{)}$. Then we define $Q^{k}_{h}(\cdot,\cdot)=\min\{(\tilde{f}^{k}_{h,\bar{\eta}})^{(1)}(\cdot,\cdot)+b^{k}_{h}(\cdot,\cdot),H\}$ and choose the greedy policy $\pi^{k}_{h}(\cdot)$ with respect to $Q^{k}_{h}$. Next, we update all $N$ normalized moments of $Q$-distribution $\psi_{1:N}\Big{(}\eta_{k}^{h}(\cdot,\cdot)\Big{)}$ and $V$-distribution $\psi_{1:N}\Big{(}\bar{\eta}_{k}^{h}(\cdot)\Big{)}$. We repeat the procedure until all the $K$ episodes are completed.
@@ 5 Theoretical Analysis
In this section, we provide the theoretical guarantees for SF-LSVI under Assumption 3.7. Applying proof techniques from Wang et al. (2020) and extending the result to a statistical functional lens, we generalize eluder dimension (Russo & Van Roy, 2013) to the vector-valued function, which has been widely used in RL literatures (Ayoub et al., 2020; Wang et al., 2020; Jin et al., 2020) to measure the complexity of learning with the function approximators.
@@
  • A state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ is $\bm{\epsilon}$-dependent on $\mathcal{Z}$ with respect to $\mathcal{F}^{N}$ if $\|f-g\|_{\mathcal{Z}}\leq\epsilon$ for any vector-valued function $f,g\in\mathcal{F}^{N}$, then $|f^{{(1)}}(s,a)-g^{{(1)}}(s,a)|\leq\epsilon$.
@@
$\left\|\tilde{f}_{h,\bar{\eta}^{\prime}}(\cdot,\cdot)-\psi_{1:N}\Big{(}(\mathcal{B}_{r(\cdot,\cdot)})_{\#}[\mathbb{P}\bar{\eta}^{\prime}](\cdot,\cdot)\Big{)}\right\|_{\mathcal{Z}_{h}^{k}}$
$\leq{c^{\prime}\left(N^{\frac{1}{2}}H\sqrt{\log(1/\delta)+\log\mathcal{N}(\mathcal{F}^{N},1/T)}\right)}$
@@
Due to the definition of Bellman unbiasedness, we remark that moment sketch has a corresponding vector-valued estimator $\phi_{\psi_{1:N}}$ as an identity and leads to a concentration results as the sampled sketches forms a martingale with respect to the filtration $\mathbb{F}^{\tau}_{h}$ induced by the history of $\{(s_{h}^{\tau},a_{h}^{\tau})\}_{\tau\in[k-1]}$ (i.e., $\mathbb{E}\Big{[}\psi_{1:N}\Big{(}(\mathcal{B}_{r_{h}})_{\#}\bar{\eta}(s^{\tau}_{h})\Big{)}\Big{|}\mathbb{F}^{\tau}_{h}\Big{]}=\psi_{1:N}\Big{(}(\mathcal{B}_{r_{h}})_{\#}[\mathbb{P}_{h}\bar{\eta}](s^{\tau}_{h},a^{\tau}_{h})\Big{)}$).
Another notable aspect in Lemma 5.3 is using normalized moments $\mathbb{E}[Z^{n}]/H^{n-1}$ instead of moments $\mathbb{E}[Z^{n}]$, as it reduces the size of the confidence region from $O(H^{N})$ to $O(\sqrt{N})$. This adjustment is akin to scaling the optimization function in multi-objective optimization to treat each objective equally, which effectively prevents the model from favoring objectives with larger scales.
@@
Under weaker structural assumptions, we show that SF-LSVI enjoys near-optimal regret bound of order $\tilde{O}(d_{E}H^{\frac{3}{2}}\sqrt{K})$, which is $\sqrt{H}$ better than the state-of-the-art distRL algorithm V-EST-LSR (Chen et al., 2024). For the linear MDP setting, we have $d_{E}=\tilde{O}(d)$ and thus SF-LSVI achieves a tight regret bound as $\tilde{O}(\sqrt{d^{2}H^{3}K})$ which matches a lower bound $\Omega(\sqrt{d^{2}H^{3}K})$ (Zhou et al., 2021). In our analysis, we highlight two main technical novelties which significantly reduces the degree of regret in distRL framework;
@@
    We refine previous lemma of Osband et al. (2019) and Wang et al. (2020) to remove the dependency of $\beta(\mathcal{F}^{N},1/\delta)$ (See Appendix D.4), ensuring that regret bound depends only on the pre-defined function class, not on the number of moment extracted.
@@ 6 Conclusions
We describe the sources of approximation error inherent in distribution-based updates and introduce a pivotal concept of Bellman unbiasedness, which enables to exactly learn the information of distribution. We also present a provably efficient online distRL algorithm, SF-LSVI, with general value function approximation. Notably, our algorithm achieves a near-optimal regret bound of $\tilde{O}(d_{E}H^{\frac{3}{2}}\sqrt{K})$, matching the tightest upper bound achieved by non-distributional framework (Zhou et al., 2021; He et al., 2023). One interesting future direction would be to reformulate the definition of regret as discrepencies in moments rather than the expected return, and to show the sample-efficiency of distRL. We hope that our work sheds some light on future research in analyzing the provable efficiency of distRL.
Acknowledgements
This work is in part supported by the National Research Foundation of Korea (NRF, RS-2024-00451435(40%), RS-2024-00413957(40%)), Institute of Information & communications Technology Planning & Evaluation (IITP, RS-2021-II212068(10%), 2021-0-00180(10%)) grant funded by the Ministry of Science and ICT (MSIT), grant-in-aid of HANHWA SYSTEMS, Institute of New Media and Communications(INMAC), and the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2025.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Notation
Appendix B Pseudocode of SF-LSVI and Technical Remarks
[…]
1:  for episode $k=1,2,\dots,K$ do
2:   Receive initial state $s_{1}^{k}$
3:   Initialize $\psi_{1:N}(\bar{\eta}_{H+1}^{k}(\cdot))\leftarrow\bm{0}^{N}$
4:   for step $h=H,H-1,\dots,1$ do
11:    $\psi_{1}\Big{(}\bar{\eta}^{k}_{h}(\cdot)\Big{)}\leftarrow V^{k}_{h}(\cdot)$,  $\psi_{2:N}\Big{(}\bar{\eta}^{k}_{h}(\cdot)\Big{)}\leftarrow\psi_{1:N}\Big{(}\eta^{k}_{h}(\cdot,\pi^{k}_{h}(\cdot))\Big{)}_{n\in[2:N]}$
12:   for $h=1,2,\dots,H$ do
13:    Take action $a^{k}_{h}\leftarrow\pi^{k}_{h}(s^{k}_{h})$
14:    Observe reward $r^{k}_{h}(s^{k}_{h},a^{k}_{h})$ and get next state $s^{k}_{h+1}$.
Remark B.1.
For an optimistic planning, we define the bonus function as the width function $b^{k}_{h}(s,a)\coloneqq w^{k}_{h}((\mathcal{F}^{N})^{k}_{h},s,a)$ where $(\mathcal{F}^{N})^{k}_{h}$ denotes a confidence region at step $h$, episode $k$. When $\mathcal{F}$ is a linear function class, the width function can be evaluated by simply computing the maximal distance of weight vector. For a general function class $\mathcal{F}$, computing the width function requires to solve a set-constrained optimization problem, which is known as NP-hard (Dann et al., 2018). However, a width function is computed simply for optimistic exploration, and approximation errors are known to have a small effect on regret (Abbasi-Yadkori et al., 2011).
Appendix C Related Work and Discussion
C.1 Technical Clarifications on Linearity Assumption in Existing Results
@@
Rowland et al. (2019) proved that quantile functional is not Bellman closed by providing a specific counterexample. However, their discussion based on counterexamples can be generalized as it assumes that the sketch Bellman operator for the quantile functional needs to be linear.
@@
Marthe et al. (2023) proposed the notion of Bellman optimizable statistical functional which redefine the Bellman update by planning with respect to statistical functionals rather than expected returns. They proved that $W_{1}$-continuous Bellman Optimizable statistical functionals are characterized by exponential utilities $\frac{1}{\lambda}\log\mathbb{E}_{Z\sim\eta}[\exp(\lambda Z)]$. However, their proof requires some technical clarification regarding the assumption that such statistical functionals are linear.
@@
C.2 Existence of Nonlinear Bellman Closed Sketch.
@@
  • $\mathcal{T}_{\psi_{\text{max}}}\Big{(}\psi_{\text{max}}(\bar{\eta}(s))\Big{)}=\max_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\Big{(}\psi_{\text{max}}\big{(}(\mathcal{B}_{r})_{\#}\bar{\eta}(s^{\prime})\big{)}\Big{)}=\max_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\Big{(}r+\psi_{\text{max}}\big{(}\bar{\eta}(s^{\prime})\big{)}\Big{)}$
  • $\mathcal{T}_{\psi_{\text{min}}}\Big{(}\psi_{\text{min}}(\bar{\eta}(s))\Big{)}=\min_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\Big{(}\psi_{\text{min}}\big{(}(\mathcal{B}_{r})_{\#}\bar{\eta}(s^{\prime})\big{)}\Big{)}=\min_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\Big{(}r+\psi_{\text{min}}\big{(}\bar{\eta}(s^{\prime})\big{)}\Big{)}$.
C.3 Non-existence of sketch Bellman operator for quantile functional
@@
Definition C.1 (mixture-consistent).
A sketch $\psi$ is mixture-consistent if for any $\nu\in[0,1]$ and any distributions $\eta_{1},\eta_{2}\in\mathscr{P}_{\psi}(\mathbb{R})$, there exists a corresponding function $h_{\psi}$ such that
$\psi(\nu\eta_{1}+(1-\nu)\eta_{2})=h_{\psi}\Big{(}\psi(\eta_{1}),\psi(\eta_{2}),\nu\Big{)}.$
@@
$\mathbb{E}_{Z\sim\nu\eta_{1}+(1-\nu)\eta_{2}}[Z^{n}\exp(\lambda Z)]=\nu\mathbb{E}_{Z\sim\eta_{1}}[Z^{n}\exp(\lambda Z)]+(1-\nu)\mathbb{E}_{Z\sim\eta_{2}}[Z^{n}\exp(\lambda Z)].$
@@
Let $\nu=\frac{1}{2}$ and $Z,Y$ be the random variables where $Z\sim\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{2}$ and $Y\sim\frac{1}{2}\delta_{k}+\frac{1}{2}\delta_{k+2}$. Then, $\text{Var}(Z)=\text{Var}(Y)=1$. While RHS is constant for any $k$, LHS is not a constant for any $k$, i.e.,
@@
$\text{Var}_{Z\sim\nu\eta_{1}+(1-\nu)\eta_{2}}[Z]$
$=\mathbb{E}_{Z\sim\nu\eta_{1}+(1-\nu)\eta_{2}}[Z^{2}]-(\mathbb{E}_{Z\sim\nu\eta_{1}+(1-\nu)\eta_{2}}[Z])^{2}$
$=\nu\mathbb{E}_{Z\sim\eta_{1}}[Z^{2}]+(1-\nu)\mathbb{E}_{Z\sim\eta_{2}}[Z^{2}]-(\nu\mathbb{E}_{Z\sim\eta_{1}}[Z]+(1-\nu)\mathbb{E}_{Z\sim\eta_{2}}[Z])^{2}$
$=\nu(\text{Var}_{Z\sim\eta_{1}}[Z]+(\mathbb{E}_{Z\sim\eta_{1}}[Z])^{2})+(1-\nu)(\text{Var}_{Z\sim\eta_{2}}[Z]+(\mathbb{E}_{Z\sim\eta_{2}}[Z])^{2})$
$\qquad-(\nu\mathbb{E}_{Z\sim\eta_{1}}[Z]+(1-\nu)\mathbb{E}_{Z\sim\eta_{2}}[Z])^{2}.$
@@
$\min_{Z\sim\nu\eta_{1}+(1-\nu)\eta_{2}}[Z]=\min(\min_{Z\sim\eta_{1}}[Z],\min_{Z\sim\eta_{2}}[Z])$
Since maximum and minimum functionals are mixture consistent, we can construct a nonlinear sketch bellman operator like the one in section C.2. This is possible because there is a nonlinear function $h_{\psi}$ that ensures the sketch is closed under mixture.
@@
Let $\nu=\frac{1}{2}$ and $Z,Y$ be the random variables where $Z\sim 0.2\delta_{0}+0.8\delta_{1}$ and $Y\sim 0.6\delta_{0}+0.4\delta_{k}$ for some $10<k<$ 1. Then $\psi_{\text{med}}(Z)=1$ and $\psi_{\text{med}}(Y)=0$. However,
@@
For a given integer $N>0$ and a quantile level $\alpha\in(0,1)$, let $\nu=\frac{1}{2}$ and a random variable $Y\sim p_{y_{0}}\delta_{0}+p_{y_{1}}\delta_{y_{1}}+\cdots+p_{y_{N}}\delta_{y_{N}}(0<y_{1}<\cdots<y_{N}<1)$ where $p_{y_{0}}>\alpha$ so that $\psi_{{\alpha}-\text{quantile}}[Y]=0$. Consider another random variable $Z\sim p_{z_{0}}\delta_{0}+p_{z_{1}}\delta_{1}$ where $p_{z_{0}}<\alpha$ so that $\psi_{{\alpha}-\text{quantile}}[Z]=1$. Then the $\alpha-$quantile of the mixture $X=\frac{Y+Z}{2}$ is
$\psi_{{\alpha}-\text{quantile}}[X]=y_{n}\text{ where }n=\min\left\{n\leq N\Big{|}\ \frac{1}{2}\sum_{n^{\prime}=0}^{n}p_{y_{n^{\prime}}}+\frac{1}{2}p_{z_{0}}>\alpha\right\}.$
Letting $p_{z_{0}}=2\alpha-\sum_{n^{\prime}=0}^{n}p_{y_{n^{\prime}}}$, we can manipulate $\psi_{{\alpha}-\text{quantile}}[X]$ to be any value of $y_{n}$. Hence, $\psi_{{\alpha}-\text{quantile}}[X]$ is a function of all possible outcomes of $Y$.
@@
Consider an MDP where initial state $s_{0}$ has no reward and transits to two state $s_{1},s_{2}$ with probability $\nu,1-\nu$ and reward distribution $\bar{\eta}_{1},\bar{\eta}_{2}$. Since $\psi$ is Bellman closed, $\psi(\bar{\eta}(s_{0}))$ is a function of $\psi(\bar{\eta}(s_{1}))$ and $\psi(\bar{\eta}(s_{2}))$, (i.e., $\psi(\bar{\eta}(s_{0}))$ = $g_{\psi}(\psi(\bar{\eta}(s_{1}))$, $\psi(\bar{\eta}(s_{2})))$ for some $g_{\psi}$). Since $\psi(\bar{\eta}(s_{0}))=\psi(\nu\bar{\eta}(s_{1})+(1-\nu)\bar{\eta}(s_{2}))$, it implies that $\psi$ is mixture-consistent. ∎
Combining the results of Lemma C.2 and Lemma C.3, we prove that a quantile sketch cannot be Bellman closed, no matter what additional sketches are provided.
Appendix D Proof